Consider the equations: f(x)=-6x-1 and g(x)=4x^2. Select the solution for (f+g)(x).
e) -6x-1+4x2
f) -24x^3-4x2
g) -24x2-1
h) -6x-1-24x2

Answer by Theo(13342)

(f+g)(x) = f(x) + g(x) = -6x-1 + 4x^2 = -6x-1+4x^2 = selection e.